# Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages

### v1

@@ 4 Experimental Settings
Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five low-resource Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^12] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE. The dataset comprises two main components:
@@
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). All translations were produced and validated by native speakers, with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
@@
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see detail in Appendix 2. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement rate of 98.5%. The corresponding Fleiss’ kappa score for semantic consistency in the DeepSeek model was 0.34, reflecting moderate inter-annotator agreement given the subjective nature of generated content.
@@
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly.
@@
Computational Resources: This study demanded significant computational resources to evaluate political bias through multiple language models and linguistic contexts. We utilized commercial API services for model access, with OpenAI API costs approximately \$287 USD for open-AI model evaluations (GPT-3.5-turbo, GPT-4, GPT-4-turbo, open-ai mini, open-ai preview) in five languages. Additional computational expenses included cloud computing resources for running open-source models and multilingual NLP pipelines for entity identification and sentiment analysis. The total computational budget, including API costs, cloud infrastructure, and model inference, was estimated at \$350 USD. Processing the complete experimental pipeline, including PCT evaluation, 440,340 headline generation, and three-method framing analysis, required approximately 120 GPU-hours on NVIDIA A100 instances.
5 Results and Analysis
5.1 Political Stance Distribution Across Languages
Figure 2 presents political stance detection results for various LLMs evaluated in Urdu, Punjabi, Balochi, Sindhi, and Pashto using the Political Compass Test (PCT). The majority of models cluster within the libertarian-left quadrant, indicating progressive economic views coupled with socially independent positions. Among them, Claude exhibits the strongest libertarian tendencies, while GPT-4-turbo scores the most pronounced leftward economic stance.
Interestingly, some models, such as OpenAI o1-mini and GPT-3.5-turbo, position in the authoritarian-right quadrant when responding in Urdu, indicating traditional or conservative leanings in this language context. Sindhi results largely reflect libertarian-left stances across models.
Overall, the PCT results reveal considerable variance in political positioning across models and languages. GPT-based models consistently shift toward authoritarian-left stances in regional languages (Urdu, Punjabi, Pashto, Balochi), whereas open-source models like Claude, Gemini, Mistral, and DeepSeek predominantly maintain libertarian-left orientations Johnson and Goldwasser (2016). This suggests that GPT models may interpret non-Western linguistic inputs as favoring stronger state control.
BERT variants generally display right-leaning tendencies irrespective of language. These findings underscore how language choice significantly influences LLMs’ expressed political biases, reinforcing the necessity for multilingual evaluations to fully capture bias dimensions.
Figure 2: Political leaning of open source and closed source models used for low-resource language shows diverse inclination across LLM
5.1.1 Cultural Adaptive Models are Less Biased
We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^13] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^14] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 3. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as:
Bias $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$
Where:
  • $x$ = position on the economic scale
  • $y$ = position on the social (authoritarian/libertarian) scale
  • $(0,0)$ = perfect neutrality (bias score = 0)
Other bias score formulations are Bias Score = |x| + |y|  [^15] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^16] Chebyshev Distance. The interpretation scale is:
Figure 3: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance and political inclination of LLMs
Model classification by bias score is measured as:
LM =$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
The closer the bias score is to zero, the more politically neutral the model is. Any deviation from the origin $(0,0)$ indicates increasing political bias. This mathematical framework clearly shows that zero represents perfect neutrality, and any movement away from the center introduces measurable bias.
The Urdu fine-tuned models demonstrated superior performance in realizing more composed and centered political positioning compared to state-of-the-art (SOTA) models. These models effectively reduced political bias and achieved improved neutrality. This outcome indicates that targeted fine-tuning on Urdu linguistic features helps models to better engage with nuanced cultural and political discourse. These models show knowledgeable and context-aware behavior in terms of political and cultural perceptions rooted specifically in Pakistan. The major implication of Fine-tuning for cultural $\rightarrow$ linguistic adaptation can integrate regional political viewpoints, producing models that are culturally aware while remaining politically neutral. This highlights the necessary balance between cultural relevance and political neutrality in AI systems.

### v2

@@ 4 Experimental Settings
Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^5] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE. The dataset comprises two main components:
@@
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages as detailed Table in Appendix 28. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). we employed three native speakers per language with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
@@
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 6. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3.
@@
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.—and were compensated fairly. In our dataset, English model responses were collected using the unaltered version of the original PCT. This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across
[…]
Computational Resources: This study required substantial computational resources to evaluate political bias across multilingual LLMs. We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines. The total budget was around \$350 USD. The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
5 Results and Analysis
5.1 Political Stance Distribution Across Languages
Figure 4 presents political stance outcomes from the Political Compass Test (PCT) across five Pakistani languages. Most LLMs cluster in the libertarian-left quadrant, reflecting progressive economic and independent social values. Claude shows the strongest libertarian stance, while GPT-4-turbo leans most economically left. Models like GPT-3.5-turbo and OpenAI o1-mini shift toward authoritarian-right in Urdu, highlighting language-specific influences. Sindhi remains consistently libertarian-left, whereas BERT variants lean right across languages. GPT models trend authoritarian-left in regional contexts, unlike open-source models which remain more liberal-libertarian. These findings underscore the need for multilingual political bias evaluation to ensure culturally equitable model behavior Johnson and Goldwasser (2016).
Figure 4: Political leaning of open source and closed source models used for Pakistani language shows diverse inclination across LLM
Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior. While English outputs tend to align with fairness and neutrality—often in the libertarian-left quadrant—responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 5. Pashto remains closest to English in political leaning. Even fine-tuned Urdu models retain or amplify these biases (Figure 13). Error analysis (Figure 16) shows English as the most stable, whereas Pakistani languages exhibit greater variance and bias. This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
Figure 5: Deterministic decoding reveals political bias shifts in LLMs across six Pakistani languages and English, exposing deep cross-linguistic bias under deterministic conditions.
Figure 6: Entity-level analysis of Urdu language outputs across LLMs. The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic. with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^6] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^7] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 13. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$. Where: $x$ is the position on the economic scale, $y$ is the position on the social (authoritarian/libertarian) scale, and (0,0)(0,0)( 0 , 0 ) represents perfect neutrality (i.e., bias score = 0). Other bias score formulations are Bias Score = |x| + |y|  [^8] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^9] Chebyshev Distance. The interpretation scale is:
Model classification by bias score is measured as:
$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
Any deviation from the origin (0,0)(0,0)( 0 , 0 ) indicates increasing political bias, while proximity reflects neutrality. Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses. This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings.
5.2 Framing Bias Analysis
Content Bias: In content bias analysis, we evaluated framing dimension and entitiy frequency analysis, which are described below.
Framing Dimension: All models exhibit varying uses of cultural identity frames when discussing religious minority issues in Pakistan as shown in Figure 12. Claude emphasizes universalist fairness, while GPT-4 integrates fairness, morality, and regulation. For the death penalty, Claude and Gemini rely on morality, whereas GPT-4 and DeepSeek favor policy frames. Shared use of constitutional framing reflects Islamic legal influence. Abstract frames lead to higher model error rates, as shown in Figure 7 Fazal (2022).
Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al. (2022). Figure 6 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing. Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
Stylistic Bias: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 14 shows some of society’s most sensitive and debated issues. Same-sex marriage rights stand out with the most polarized coverage—yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse. Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone. On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment. When it comes to government performance, coverage tends to be both critical and sharply divided—reflecting growing public frustration as shown in Figure 17. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression. For more analysis and results see Appendix C.
Figure 7: Error rates across political discourse dimensions reveal how confidently models handle different types of content.
6 Findings
6.1 Model-Level Bias Interpretation
Figure  8 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. To further interpret how political bias manifests in model behaviour, we conducted a detailed, multidimensional analysis of the DeepSeek model’s visual and quantitative breakdowns, which are provided in the Appendix Figure 21.
Figure 8: Lexical Polarity Rate (LPR) comparison across 11 sociopolitical topics for LLMs.
6.2 Cross-Cultural model insights
Our analysis reveals distinct model behaviors across cultural-linguistic contexts. GPT models show high cultural sensitivity, shifting from libertarian-left in English to authoritarian-left in Pakistani languages (Figure 4). Claude remains consistently libertarian with progressive sentiment. Open-source models vary—Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts. These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance—something current models often fail to achieve.
6.3 Multilingual Influence on Political Bias
The clustering plots of political stance detection across languages reveal that LLMs exhibit biased behavior influenced by political and cultural context. This variation highlights imbalances in training data, with English-dominant models adapting differently in low-resource languages. Figure 18 shows consistent support for topics like Education but sharp divergence on Language Policy. These findings suggest that multilingualism can generate language-specific political personalities, raising concerns that global LLM deployment may unintentionally reinforce cultural biases depending on the language used.
7 Conclusion
This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi. We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis—capturing both what the model says and how it says it. Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior. By bridging the Political Compass Test (PCT) with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs. This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
Limitations
This study has several methodological and scope-related limitations. Our analysis is restricted to five Pakistani languages, omitting other regional or global languages that may exhibit different ideological patterns. We also exclude high-resource languages like French or Arabic, which could have served as cultural counterfactuals to help distinguish between linguistic and cultural influences in LLM pretraining. While we adapt the Political Compass Test (PCT), its original design is rooted in Western political thought and may not fully capture culturally specific structures like Islamic jurisprudence or tribal governance. Additionally, our sentiment and entity analysis tools are primarily trained on English, potentially reducing accuracy and missing localized expressions. We do not explore prompt-induced framing bias, which may affect stance independently of model ideology. Finally, the static nature of training data limits our ability to assess real-time political shifts. Future work should explore culturally sensitive bias mitigation for low-resource settings.
Ethical Statement
Appendix A Appendix
A.1 Dataset Contribution
Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five languages shown in the Table 28. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  9. The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
A.2 Language Translation Procedure
For each of the five target languages, prompts were translated from English by a team of three bilingual native speakers. This was followed by model generation in the respective language, forming a two-step pipeline. While multi-translator input reduced individual bias, we acknowledge the absence of back-translation or inter-annotator agreement as a limitation that may introduce subtle framing or cultural interpretation bias in multilingual political contexts.
A.3 Translation Assignment Procedure
Each statement was independently translated into the target language by three bilingual native speakers. Translations were then compared, and a final version was selected through majority agreement or collaborative consensus when needed. This approach ensured semantic accuracy while reducing individual translator bias. For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES. The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement. The interpretations are based on standard kappa ranges, that is:
[…]
The Fleiss’ kappa statistic is calculated as:
$\kappa=\frac{P-P_{e}}{1-P_{e}}$
The significance of calculating the kappa score for Pakistani languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages. Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
@@ A.3 Translation Assignment Procedure
Appendix B Technical details
B.1 Stance Detection vs. Our Ideological Mapping
Standard stance detection typically involves classifying whether a text expresses a pro, con, or neutral position toward a specific target or claim Mohammad et al. (2016). These approaches focus on binary or ternary stance concerning an explicit target, often in single-turn texts such as tweets. In contrast, our methodology uses an adapted Political Compass Test to infer a model’s position in a two-dimensional ideological space (economic and social axes). Rather than target-specific classification, we aggregate stance scores across 62 political statements to construct a holistic ideological profile per model and language. This offers a structured lens into political bias beyond isolated stance decisions. Moreover, we complement this scalar stance mapping with rhetorical framing analysis to examine how ideological leanings are expressed stylistically and narratively—going beyond traditional stance detection’s limited focus on polarity or agreement.
Stance detection is a widely studied task in NLP Gorrell et al. (2019), typically formulated as predicting whether a speaker is in favor or against a known target or topic. These tasks are often applied to tweets, debates, or news articles. In contrast, our use of the Political Compass Test allows for continuous stance scoring across a spectrum of ideologically salient statements, enabling two-dimensional mapping of model behavior. This richer representation is particularly important for analyzing latent political bias in generative LLMs, where stance is not tied to a single topic but emerges across diverse ideological domains.
B.2 Prompting Strategy and Reproducibility
We adopt a zero-shot prompting strategy for all LLM generations to avoid bias from exemplars. For each of the 62 PCT statements, models are asked to respond in the specified language using a Likert-style agreement format. For framing analysis, we generate 1,000 headlines per stance–topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 21). Frame classification is conducted using GPT-3.5-turbo with a fixed schema based on Boydstun’s taxonomy. Named entity and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa. All code and prompts are shared publicly for reproducibility.
B.3 Model Details
Table 2 provides an overview of the language models used in our bias evaluation. It includes both closed-source and open-source models, detailing their type, estimated parameter sizes, and architectures. Hyperlinked model names direct to official documentation or repositories, enabling transparency and reproducibility for further comparative analysis.
Model Architectures and Tuning Details Most of the models we evaluate, including GPT-4, Claude, Gemini, and DeepSeek, are not raw pretrained models but represent fully developed systems with instruction tuning and safety alignment, often including RLHF. Our focus is on assessing bias as it appears in real-world, user-facing outputs. We include both decoder-based models (e.g., DeepSeek-Chat) and encoder-based models (e.g., BERT, RoBERTa) to explore architectural effects. While decoder models are prompted generatively, encoder models are probed using classification on the same inputs. We acknowledge that this mix, along with post-training layers, may influence results and that further work is needed to isolate these factors.
Table 2: Overview of Language Models Used in Bias Evaluation
Model Name Type Parameters Architecture
GPT-3.5-turbo Closed-source $\sim$175B (est.) Decoder
GPT-4-turbo Closed-source $\sim$1.8T (est.) Decoder
GPT-4 Closed-source $\sim$1.8T (est.) Decoder
GPT-4o Closed-source $\sim$1.8T (est.) Decoder
OpenAI o1-mini Closed-source Unknown Decoder
OpenAI o1-preview Closed-source Unknown Decoder
Claude-3-Haiku-202403 Closed-source $\sim$13B (est.) Decoder
Gemini-1.5-Pro Closed-source Unknown Decoder
Mistral-7B-Instruct-v0.2 Open-source 7B Decoder
DeepSeek-Chat Open-source 7B Decoder
XLM-RoBERTa-large Open-source 550M Encoder
BERT-large Open-source 340M Encoder
XLM-RoBERTa-base Open-source 270M Encoder
B.4 On Bias and Limitations of mDeBERTa
While mDeBERTa-v3-base-mnli-xnli offers strong cross-lingual performance for zero-shot stance classification, we acknowledge that it may carry latent biases inherited from its training on the XNLI corpus. The XNLI dataset is primarily derived from translations of English data and may over-represent high-resource languages and Western discourse norms. This could influence how disagreement or ambiguity is expressed in lower-resource languages like Urdu or Balochi. Although mDeBERTa outperformed alternatives such as XLM-R and mBERT in pilot tests for our target languages, we note that future work should explore culturally fine-tuned models or adversarial probing to better surface language-specific classification bias.
Role of mDeBERTa in Stance Classification. To perform stance classification over multilingual PCT responses, we utilized mDeBERTa-v3-base, a multilingual variant of DeBERTa pretrained on XLM-R corpora, which offers enhanced cross-lingual representation capabilities. We selected mDeBERTa over alternatives such as XLM-RoBERTa and mBERT due to its superior performance in zero-shot stance and sentiment classification benchmarks, especially for underrepresented languages. Its disentangled attention mechanism and language-agnostic pretraining make it a suitable choice for capturing ideological nuance across the five Pakistani languages evaluated. We fine-tuned mDeBERTa on translated PCT examples and constrained the output to four stance labels (Strongly Agree, Agree, Disagree, Strongly Disagree). While mDeBERTa performs competitively, we acknowledge potential limitations from English-centric pretraining that may introduce biases or misalignments in culturally specific contexts, which we mitigate through triple-verified human translations and language-specific prompt tuning. Future work can explore culturally grounded multilingual encoders tailored to South Asian political discourse.
B.5 Stance Classification Details
We use the mDeBERTa-v3-base-mnli-xnli model from HuggingFace’s Transformers library as a zero-shot classifier to assign stance labels. The model is prompted with concatenated input: the PCT statement and model response. It returns softmax confidence scores across four labels: Strongly Agree, Agree, Disagree, and Strongly Disagree. The label with the highest confidence is selected as the predicted stance. We compute a numeric stance score by multiplying the winning label’s score by ±5 or ±10. For example, a confidence of 0.86 on Strongly Disagree yields a stance score of –8.6. This scoring method provides interpretable agreement strength in both ordinal and continuous forms.mDeBERTa-v3-base-mnli-xnli was chosen for stance classification because of its zero-shot multilingual capabilities and language coverage, outperforming alternatives like XLM-R or monolingual classifiers in cross-lingual consistency.
B.6 Standard PCT Automation
We follow the original two-dimensional structure of the Political Compass Test, consisting of the economic (Seco) and social (Ssoc) axes. Rather than replicate its internal scoring algorithm, we simulate user input by mapping model-generated stance scores to the 4-option Likert scale and input them programmatically into the official PCT interface using Selenium. This yields authentic quadrant-level coordinates directly from the source.
B.7 Language Focus and Fine-tuning Scope.
In this study, we did not conduct full model fine-tuning due to resource constraints and instead relied on zero-shot and instruction-tuned responses from existing large language models (LLMs) across multiple languages. For controlled probing in low-resource settings, we strategically focused on Urdu to examine how political and economic bias manifests in culturally grounded contexts. Urdu was selected based on its linguistic richness, wider resource availability, and its status as the mother tongue of Pakistan.
B.8 Framing Setup
We acknowledge that using a single model response per prompt may introduce sampling variance in frame predictions. Incorporating majority voting across multiple generations, or ensembling across different models, could improve the robustness of frame assignment and reduce random variability. We leave this to future work due to API cost constraints.
B.9 Sentiment Classifier
We employ XLM-RoBERTa (XLM-R) for downstream tasks such as multilingual frame classification due to its robust performance across 100 languages, including low-resource ones like Urdu, Punjabi, and Pashto. Trained on CommonCrawl data in a self-supervised manner, XLM-R provides strong cross-lingual generalization, making it well-suited for tasks where labeled data is scarce or unavailable in the target language. Compared to alternatives like mBERT, XLM-R achieves superior results in cross-lingual transfer, particularly for sentence-level classification tasks, while maintaining consistency across diverse scripts. Its architecture also allows effective integration with frame-tagging pipelines in our zero-shot or few-shot evaluations.
Appendix C Detailed Results
C.1 Ideological Consistency of PCT Responses
Figure 10 illustrates the distribution of model responses to Political Compass Test (PCT) prompts projected along the economic (Seco) and social (Ssoc) axes, with bootstrapped 95% confidence intervals shown as translucent error bars. Each dot corresponds to a single prompt, and its position reflects the average ideological stance expressed by the model. The plot reveals a strong diagonal clustering from the lower-left to the upper-right quadrant, indicating a high correlation between the model’s economic and social leanings. Despite variation introduced through multilingual translations, the narrow spread of error bars for most points suggests stable and consistent model behavior. A few prompts with wider intervals reflect ideologically ambiguous or culturally sensitive content. This visualization offers a fine-grained and interpretable view of model ideology, moving beyond discrete stance labels and enabling deeper insight into alignment patterns across ideological dimensions.
C.2 Ideological Leanings of LLMs Across Political Topics
The heat-map in Figure  18 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change. Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances. Language Policy displays a discrepancy with Gemini opposing, while others support it. Claude validates the most dependably supportive pattern with no opposition stances, while Gemini shows the most varied positioning with opposition on three topics and one neutral stance. Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
Figure 10: Mean stance positions on economic (Seco) and social (Ssoc) axes for PCT prompts, with 95% bootstrapped confidence intervals ($n=10$). The diagonal pattern indicates a strong correlation between ideological dimensions.
Figure 11: Stance comparison Religious Minorities Rights